# Nearest Neighbour Score Estimators for Diffusion Generative Models

### v1

@@ C.2 Uniform SNIS Trace of Covariance
[…]
@@ D.2 Additional Denoiser Images
Figure 8 show additional KNN posterior mean estimates for a variety of source images and noise levels.
@@ E.1 Consistency Model Training Configurations
To train the consistency models reported in Table 1, we follow the configuration outlined by Song & Dhariwal (2023). For the reader’s convenience we reiterate that configuration here. We use the NCSN++ architecture (Song et al., 2021) for all models. Each model was trained using the improved consistency training algorithm, with the Pseudo-Huber loss function (Charbonnier et al., 1997) with $c=0.03$. We weight the consistency matching loss with the weighting function $\lambda(t)=\frac{1}{\sigma_{t}-\sigma{t-1}}$. Noise levels are sampled from a log-normal distribution with $P_{mean}=-1.1$ and $P_{std}=2.0$. We follow an exponential discretzation schedule, starting with $s_{0}=10$ discretization steps, doubling every 50,000 training iterations to a final value of $s_{1}=1280$. We do not use exponential averaging on the teacher network, instead just applying a stopgrad to the student network weights. Hyperparameters for all runs are reported in Table 2

### v2

@@ C.2 Uniform SNIS Trace of Covariance
Let $\hat{\mu}_{\text{U}}$ be a SNIS posterior mean estimator with $q(\mathbf{x}^{(i)})=\frac{1}{N}\ \forall\mathbf{x}^{(i)}\in D$ then,
$n\cdot\text{Tr}\left(\text{Cov}\left(\hat{\mu}_{\text{U}}\right)\right)=N\cdot\sum_{i=1}^{N}p_{t}(\mathbf{x}^{(i)}|\mathbf{z})^{2}(\mathbf{x}^{(i)}-\mathbb{E}\left[\mathbf{x}|\mathbf{z},t\right])^{\circ 2}.$ (78)
@@ D.2 Additional Denoiser Images
Figure 8 shows additional KNN posterior mean estimates for a variety of source images and noise levels.
@@ D.3 CIFAR-10 Consistency Model Samples
D.4 Diffusion Models
Although we demonstrate our method’s effectiveness for consistency model training in Section 5, our method is not specifically tailored to consistency models, and may be used during model training whenever a empirical score estimator is required. In particular, our estimator can be directly applied to train diffusion models.
To evaluate the effect of our estimator on diffusion model training, we train and evaluate unconditional EDM (Karras et al., 2022) models on CIFAR-10 (Krizhevsky et al., 2009) using varying score estimators. Specifically, we train models using the standard single-sample Monte Carlo estimator, the STF estimator (Xu et al., 2023), and our nearest neighbour estimator. We retrain EDM and STF from their official repositories, utilizing the DDPM++ architecture for all models. We report the best FID and IS on 50,000 samples across all model checkpoints in Table 2
Table 2: Unconditional EDM CIFAR-10 sample quality with varying score estimators.
[…]
EDM (Karras et al., 2022) 2.01 9.76
STF (Xu et al., 2023) 1.94 9.81
EDM + KNN (ours) 1.96 9.87
From Table 2, we see that both our method and STF improve the quality of generated samples over EDM. Compared to STF, our method has similar FID and better Inception Score.
D.5 Score Estimation on Additional Datasets
We apply our approach to two additional datasets to evaluate the efficacy of our method on larger and higher dimensionality data. Figure 10 shows the performance of our score estimator on CelebA 64x64 (Liu et al., 2015), while Figure 11 demonstrates the score estimation error of out method on Imagenet 64x64 (Deng et al., 2009).
Figure 10: Comparison of various score estimators on unconditional CelebA 64x64
Figure 11: Comparison of various score estimators on unconditional Imagenet 64x64
Applying our method to higher dimensional and larger datasets presents two challenges. Firstly, to perform fast nearest neighbours search, we generate an index of the entire dataset which is stored in memory. Larger images and larger datasets therefore present memory concerns. Secondly, larger datasets increase the runtime of nearest neighbour search, which especially apparent on large datasets such as Imagenet.
To address these challenges, we apply two strategies. To reduce the memory overhead of our index, we quantize the data to 8 bit precision, reducing memory overhead by a factor of four. This has a minimal affect on our estimator, as pixel intensities are generally represented as 8 bit values. To improve the runtime of nearest neighbour search, we utilize an inverted file index (IVF) which partitions the data into $c$ clusters and performs nearest neighbour search over only the $p$ closest partitions. Both $c$ and $p$ can be tuned to trade off runtime versus accuracy. For CelebA 64x64 we utilized $c=1788$ and $p=25$ while for Imagenet, we used $c=4527$ and $p=100$.
In our work, we assume that the nearest neighbour search returns the $k$ most probable elements of the posterior distribution. However, by switching to an approximate nearest neighbour search, this may no longer be true. We evaluate the impact of using approximate nearest neighbour search on estimator performance in Figure 12.
Figure 12: Comparison of approximate and exact nearest neighbour score estimators on CelebA 64x64
From Figure 12, we can see that approximate nearest neighbour introduces some additional bias to the posterior mean estimation, however the bias is still less that STF for the majority of noise levels. We expect bias could be reduced by increasing the number of samples $n$ or the number of search partitions $p$.
@@ E.1 Consistency Model Training Configurations
To train the consistency models reported in Table 1, we follow the configuration outlined by Song & Dhariwal (2024). For the reader’s convenience we reiterate that configuration here. We use the NCSN++ architecture (Song et al., 2021) for all models. Each model was trained using the improved consistency training algorithm, with the Pseudo-Huber loss function (Charbonnier et al., 1997) with $c=0.03$. We weight the consistency matching loss with the weighting function $\lambda(t)=\frac{1}{\sigma_{t}-\sigma_{t-1}}$. Noise levels are sampled from a log-normal distribution with $P_{mean}=-1.1$ and $P_{std}=2.0$. We follow an exponential discretzation schedule, starting with $s_{0}=10$ discretization steps, doubling every 50,000 training iterations to a final value of $s_{1}=1280$. We do not use exponential averaging on the teacher network, instead just applying a stopgrad to the student network weights. Hyperparameters for all runs are reported in Table 3